# On functoriality and the tensor product property in noncommutative tensor-triangular geometry

### v1

@@ 4. Functoriality and universality of the complete prime spectrum
We have that $F$ factors as a surjective functor $F:\mathscr{K}\to F(\mathscr{K})$ composed with the inclusion $\iota:F(\mathscr{K})\subseteq\mathscr{L}$. It follows that $\operatorname{Spc}(\iota)$ is well-defined if and only if every prime $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$ satisfies that $\mathscr{Q}\cap F(\mathscr{K})$ is prime in $F(\mathscr{K})$, since the inverse image of $\iota$ is simply intersection. Now, if that property is satisfied, it follows that the composition $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map $\operatorname{Spc}(F)$ on Balmer spectra by Proposition 4.2. Conversely, if the property is not satisfied, then there exists some prime $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$ for which there exist thick $\otimes$-ideals $\mathscr{I},\mathscr{J}$ of $F(\mathscr{K})$ for which $\mathscr{I}\otimes\mathscr{J}\subseteq\mathscr{Q}\cap F(\mathscr{K})$ but neither $\mathscr{I}\subseteq\mathscr{Q}\cap F(\mathscr{K})$ nor $\mathscr{J}\subseteq\mathscr{Q}\cap F(\mathscr{K})$. Pulling back, it follows that $F^{-1}(\mathscr{I})\otimes F^{-1}(\mathscr{J})\subseteq F^{-1}(\mathscr{Q})$ but neither $F^{-1}(\mathscr{I})\subseteq F^{-1}(\mathscr{Q})$ nor $F^{-1}(\mathscr{J})\subseteq F^{-1}(\mathscr{Q})$. Since $F^{-1}(\mathscr{I})$ and $F^{-1}(\mathscr{J})$ are thick $\otimes$-ideals of $\mathscr{K}$, we conclude that $F^{-1}(\mathscr{Q})$ is nonprime, as desired. Continuity follows the same as before. ∎
We now turn to the complete prime spectrum, $\operatorname{Spc}^{cp}$. The next lemma is standard, key for proving universality, and follows essentially the same as the symmetric case.
@@
Now we prove, as promised, that $\operatorname{Spc}^{cp}(\mathscr{K})$ is universal: it is the final object of $\mathscr{S}(\mathscr{K})$.
@@
The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final multiplicative support data on $\mathscr{K}$. That is, if it is nonempty, $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is a multiplicative support data, and for any other multiplicative support data $(Y,\sigma)$ on $\mathscr{K}$, there exists a unique continuous map $f\colon Y\to\operatorname{Spc}^{cp}(\mathscr{K})$ such that $\sigma(x)=f^{-1}(\operatorname{supp}^{cp}(x))$. Explicitly, the map is defined by
@@
On the other hand, if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$, then $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp})$ is the only multiplicative support datum on $\mathscr{K}$, therefore it is trivially the final object of $\mathscr{S}(\mathscr{K})$.
@@
If every prime ideal of $\mathscr{K}$ is completely prime, then the inclusion $\operatorname{Spc}^{cp}(\mathscr{K})\hookrightarrow\operatorname{Spc}(\mathscr{K})$ is the identity, hence $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ is multiplicative. Conversely, if $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ has the tensor product property, then the unique map map $f:\operatorname{Spc}(\mathscr{K})\to\operatorname{Spc}^{cp}(\mathscr{K})$ defined by $f(\mathscr{P})=\{a\in\mathscr{K}\mid\mathscr{P}\not\in\operatorname{supp}(a)\}$ is clearly the identity. Thus $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})=(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$, so $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ has the tensor product property. ∎
@@
Finally, we show $\mathscr{S}^{cp}(-)$, $\operatorname{Spc}^{cp}$ is fully functorial, as promised.
@@
First we well-definedness. It is easy to check that $F^{-1}(\mathscr{P})$ is a thick $\otimes$-ideal. For primality, suppose $a\otimes b\in F^{-1}(\mathscr{P})$, then $F(a)\otimes F(b)\in\mathscr{P}$. By primality, either $F(a)$ or $F(b)$ belongs to $\mathscr{P}$, hence $a$ or $b$ belongs to $F^{-1}(\mathscr{P})$ by definition. Note that this implies that if $\operatorname{Spc}^{cp}(\mathscr{L})$ is nonempty, then $\operatorname{Spc}^{cp}(\mathscr{K})$ necessarily is as well. On the other hand, if $\operatorname{Spc}^{cp}(\mathscr{L})$ is empty, this poses no problems since the empty set is the initial object in the category of topological spaces.
@@
$\operatorname{supp}^{cp}_{\mathscr{L}}(F(x))=\{Q\in\operatorname{Spc}^{cp}(\mathscr{L})\mid F(x)\neq Q\}.$
@@
$(\operatorname{Spc}F)^{-1}\big{(}\operatorname{supp}^{cp}_{\mathscr{K}}(x)\big{)}=\{Q\in\operatorname{Spc}^{cp}(\mathscr{L})\mid x\not\in(\operatorname{Spc}(F))(\mathscr{Q})\}.$
@@
Since the compliments of supports form a basis of the topology of the complete prime spectrum, it follows that $\operatorname{Spc}(F)$ is continuous. Finally, $\operatorname{Spc}^{cp}(G\circ F)=\operatorname{Spc}^{cp}(F)\circ\operatorname{Spc}^{cp}(G)$ is easy to verify. ∎
Functoriality provides an easy check for when $\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$, the existence of a “fiber functor” to a more well-understood category. Though in the case of the stable category of a monoidal category admitting a fiber functor to $\mathbf{Vec}$, this is not much help since $\operatorname{stab}(\mathbf{Vec})\cong\{0\}$.
@@ 5. Radical thick tensor-ideals
A pertinent question with an unknown answer is whether every radical thick $\otimes$-ideal of $\mathscr{K}$ is an intersection of completely prime thick $\otimes$-ideals. On the other hand, an intersection of prime thick $\otimes$-ideals is not necessarily radical, since not all prime thick $\otimes$-ideals are radical. See Example 7.6.

### v2

@@ 4. Partial functoriality of the noncommutative Balmer spectrum
We show the statement for the noncommutative Balmer spectrum; the latter statement follows analogously. Well-definedness of $\varphi:=\operatorname{Spc}(q)$ is guaranteed by the previous proposition. Set $V:=\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid\mathscr{I}\subseteq\mathscr{P}\}$. For any $\mathscr{Q}\in\operatorname{Spc}(\mathscr{K}/\mathscr{I})$, we have $\varphi(\mathscr{Q})=q^{-1}(\mathscr{Q})\supset q^{-1}(0)=\mathscr{I}$, or in other words, $\operatorname{im}(\varphi)\subseteq V$. Proposition 5.7 asserts $\varphi$ is injective, since $q$ is surjective. Conversely. if $\mathscr{P}\in\operatorname{Spc}(\mathscr{K})$ contains $\mathscr{I}$, then it is straightforward to see $q(\mathscr{P})$ is a prime thick $\otimes$-ideal, and $q^{-1}(q(\mathscr{P}))=\mathscr{P}$, so $\varphi$ induces a continuous bijection $\operatorname{Spc}(\mathscr{K}/\mathscr{I})\to V$.
Now, let $x\in\mathscr{K}$ and $\mathscr{P}\in V$. We have $x\in\mathscr{P}$ if and only if $x\in q(\mathscr{P})$ (with $x$ regarded as an object of $\mathscr{K}/\mathscr{I}$), so $\varphi(\operatorname{supp}_{\mathscr{K}/\mathscr{I}}(x))=\operatorname{supp}_{\mathscr{K}}(x)\cap V$, hence $\varphi:\operatorname{Spc}(\mathscr{K}/\mathscr{I})\to V$ is a closed map, thus a homeomorphism as desired. ∎
We can in fact give a criterion for when an exact $\otimes$-functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on noncommutative Balmer spectra.
Theorem 4.4.
Let $F\colon\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor, and denote its image by $F(\mathscr{K})\subseteq\mathscr{L}$. Then $F$ induces a continuous map on Balmer spectra $\varphi:=\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ given by $\mathscr{Q}\mapsto F^{-1}(\mathscr{Q})$ if and only if for all prime thick $\otimes$-ideals $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, the intersection $F(\mathscr{K})\cap\mathscr{Q}$ satisfies the primality property,
$\text{For all $a,b\in F(\mathscr{K})$, }a\otimes F(\mathscr{K})\otimes b\subseteq F(\mathscr{K})\cap\mathscr{Q}\text{ implies }a\in\mathscr{Q}\text{ or }b\in\mathscr{Q}.$
Proof.
It is routine to verify that for every thick $\otimes$-ideal $\mathscr{J}\subseteq\mathscr{L}$, $F^{-1}(\mathscr{J})$ is a thick $\otimes$-ideal of $\mathscr{K}$. Therefore, for $\operatorname{Spc}(F)$ is well-defined if and only if the primality property is preserved for all primes $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$. We have that $F$, as a monoidal (but not necessarily triangulated, since $F(\mathscr{K})$ may not be triangulated if $F$ is not full) functor factors as the surjective monoidal functor $F:\mathscr{K}\to F(\mathscr{K})$ composed with the inclusion $\iota:F(\mathscr{K})\subseteq\mathscr{L}$. The same argument as Proposition 4.2 demonstrates that for every $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, $F^{-1}(F(\mathscr{K})\cap\mathscr{Q})\subseteq\mathscr{K}$ satisfies the tensor product property by surjectivity of $F:\mathscr{K}\to F(\mathscr{K}).$ We turn to the inclusion $\iota$.
If $\iota^{-1}$ preserves the primality property for all intersections of primes $\mathscr{Q}\cap F(\mathscr{K})$ with $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, it follows that the composition $(\iota\circ F)^{-1}=F^{-1}\circ\iota^{-1}$ preserves the property as well, so $\operatorname{Spc}(F)$ is well-defined. Conversely if the property is not satisfied, then there exist $a,b\in F(\mathscr{K})$ such that $a\otimes F(\mathscr{K})\otimes b\in F(\mathscr{K})\cap\mathscr{Q}$ but $a,b\not\in\mathscr{Q}$. Therefore, there exist $a^{\prime},b^{\prime}\in\mathscr{K}$ with $F(a^{\prime})=a$, $F(b^{\prime})=b$ such that $a^{\prime}\otimes\mathscr{K}\otimes b^{\prime}\subseteq F^{-1}(\mathscr{Q})=F^{-1}(\mathscr{Q}\cap F(\mathscr{K}))$, but $a^{\prime},b^{\prime}\not\in F^{-1}(\mathscr{Q})$, hence $\operatorname{Spc}(F)$ is not well-defined. Continuity follows the same as in Proposition 4.2. ∎
We obtain a standard tensor-triangular geometric result for noncommutative tensor-triangular geometry: idempotent completion does not affect the Balmer spectrum. This is a noncommutative version of [Bal05, Proposition 3.13].
Theorem 4.5.
Let $\mathscr{L}$ be a monoidal triangulated category and let $\mathscr{K}\subseteq\mathscr{L}$ be a full monoidal triangulated subcategory with the same unit and which is cofinal, i.e. for any object $x\in\mathscr{L}$, there exists $y\in\mathscr{L}$ such that $x\oplus y\in\mathscr{K}$. Then the map $\mathscr{Q}\mapsto\mathscr{Q}\cap\mathscr{K}$ defines a homeomorphism $\operatorname{Spc}(\mathscr{L})\xrightarrow{\cong}\operatorname{Spc}(\mathscr{K})$. Moreover, this restricts to a homeomorphism $\operatorname{Spc}^{cp}(\mathscr{L})\xrightarrow{\cong}\operatorname{Spc}^{cp}(\mathscr{K})$.
Proof.
First, note we may replace $\mathscr{K}$ with its isomorphic-closure, so we can assume $\mathscr{K}$ is replete. The map $\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$, if it is well-defined, is simply $\operatorname{Spc}(\iota)$, where $\iota:\mathscr{K}\hookrightarrow\mathscr{L}$ denotes inclusion. To see $\operatorname{Spc}(\iota)$ is well-defined, for any $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, $x\otimes(z\oplus z^{\prime})\otimes y\in\mathscr{Q}$ implies $x\otimes z\otimes y\in\mathscr{Q}$ and $x\otimes z^{\prime}\otimes y\in\mathscr{Q}$ since $\mathscr{Q}$ is thick. Therefore by cofinality, for any $x,y\in\mathscr{K}$, $x\otimes\mathscr{K}\otimes y\subseteq\mathscr{Q}$ implies $x\otimes\mathscr{L}\otimes y\subseteq\mathscr{Q}$ implies $x\in\mathscr{Q}$ or $y\in\mathscr{Q}$, so $\mathscr{Q}\cap\mathscr{K}$ is indeed a prime thick $\otimes$-ideal of $\mathscr{K}$, and $\operatorname{Spc}(\iota)$ is indeed a well-defined map on Balmer spectra.
Now, recall that for any $x\in\mathscr{L}$, we have $x\oplus\Sigma x\in\mathscr{K}$ (c.f. the proof of [Bal05, Proposition 3.13]). Moreover, if $x\oplus x^{\prime}$ belongs to a triangulated subcategory of $\mathscr{K}$ (e.g. a prime), then so does $x\oplus\Sigma x$. Therefore, given a prime $\mathscr{P}\in\operatorname{Spc}(\mathscr{K})$, we have an equality of full subcategories
$\widetilde{\mathscr{P}}:=\{x\in\mathscr{L}\mid x\oplus\Sigma x\in\mathscr{P}\}=\{x\in\mathscr{L}\mid\exists x^{\prime}\in\mathscr{L}\text{ such that }x\oplus x^{\prime}\in\mathscr{P}\}.$
We claim $\widetilde{\mathscr{P}}$ is a prime of $\mathscr{L}$. It is routine to verify $\widetilde{\mathscr{P}}$ is a thick $\otimes$-ideal. To verify it is prime, suppose $x\otimes z\otimes y\in\widetilde{\mathscr{P}}$ for all $z\in\mathscr{L}$, and suppose $x\not\in\widetilde{\mathscr{P}}.$ Set $w:=x\oplus\Sigma x$, then by construction, $w\in\mathscr{K}\setminus\mathscr{P}$, but
$w\otimes z\otimes y\cong(x\otimes z\otimes y)\oplus\Sigma(x\otimes z\otimes y)\in\mathscr{P}$
for all $z\in\mathscr{K}$, therefore
$w\otimes z\otimes(y\oplus\Sigma y)\cong(w\otimes z\otimes y)\oplus\Sigma(w\otimes z\otimes y)\in\mathscr{P}$
for all $z\in\mathscr{K}$, hence $y\oplus\Sigma y\in\mathscr{P}$ by primality of $\mathscr{P}$. Thus, $y\in\widetilde{\mathscr{P}}$, as desired.
It follows by $\mathscr{P}$ thick that $\widetilde{\mathscr{P}}\cap\mathscr{K}=\mathscr{P}$, so the assignment $\mathscr{P}\mapsto\widetilde{\mathscr{P}}$ is a right inverse to $\operatorname{Spc}(\iota)$. On the other hand, if $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, we claim $\mathscr{Q}=\widetilde{\mathscr{P}}$ where $\mathscr{P}:=\mathscr{Q}\cap\mathscr{K}$. Indeed, if $x\in\mathscr{Q}$, then $x\oplus\Sigma x\in\mathscr{K}$, hence $x\oplus\Sigma x\in\mathscr{Q}$, so $\mathscr{Q}\subseteq\widetilde{\mathscr{P}}$. Conversely, $\widetilde{\mathscr{P}}\subseteq\mathscr{Q}$ by thickness of $\mathscr{Q}$. So $\operatorname{Spc}(\iota)$ is a continuous bijection with inverse given by $\mathscr{P}\mapsto\widetilde{\mathscr{P}}$.
Now, for any $x\in\mathscr{L}$, and $\mathscr{P}\in\operatorname{Spc}(\mathscr{K})$, we have $x\in\widetilde{\mathscr{P}}$ if and only if $x\oplus\Sigma x\in\widetilde{\mathscr{P}}$ if and only if $x\oplus\Sigma x\in\mathscr{P}$. Therefore, the image of $\operatorname{Spc}(\iota)$ of $\operatorname{supp}_{\mathscr{L}}(x)$ is $\operatorname{supp}_{\mathscr{K}}(x\oplus\Sigma x),$ so $\operatorname{Spc}(\iota)$ is a closed map, and thus a homeomorphism, as desired.
For the final statement, one needs to verify that $\operatorname{Spc}(\iota)$ and the assignment $\mathscr{P}\mapsto\widetilde{\mathscr{P}}$ for $\mathscr{P}\in\operatorname{Spc}(\mathscr{K})$ preserve complete primality. The former follows from functoriality of $\operatorname{Spc}^{cp}$, proven in Theorem 5.5. For the latter, suppose $x\otimes y\in\widetilde{\mathscr{P}}$ and $x\not\in\widetilde{\mathscr{P}}$. Let $z:=x\oplus\Sigma x$, then we have $z\in\mathscr{K}\setminus\mathscr{P}$ and $z\otimes y\cong(x\otimes y)\oplus\Sigma(x\otimes y)\in\mathscr{P}$. Therefore, $z\otimes(y\oplus\Sigma y)\cong(z\otimes y)\oplus\Sigma(z\otimes y)\in\mathscr{P}$, so $y\oplus\Sigma y\in\mathscr{P}$, hence $y\in\widetilde{\mathscr{P}}$. ∎
Corollary 4.6.
Let $\mathscr{K}^{\natural}$ denote the idempotent completion of $\mathscr{K}$, and let $\iota:\mathscr{K}\hookrightarrow\mathscr{K}^{\natural}$ denote the canonical inclusion. The induced map on the noncommutative Balmer spectrum $\operatorname{Spc}(\iota):\operatorname{Spc}(\mathscr{K}^{\natural})\to\operatorname{Spc}(\mathscr{K})$ is a well-defined homeomorphism, which restricts to a homeomorphism $\operatorname{Spc}^{cp}(\iota):\operatorname{Spc}^{cp}(\mathscr{K}^{\natural})\to\operatorname{Spc}^{cp}(\mathscr{K})$.
5. Functoriality and universality of the complete prime spectrum
@@
Now we prove, as promised, that $\operatorname{Spc}^{cp}(\mathscr{K})$ is universal: it is the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$.
@@
The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final multiplicative support data on $\mathscr{K}$, i.e. the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$. That is, if it is nonempty, $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is a multiplicative support data, and for any other multiplicative support data $(Y,\sigma)$ on $\mathscr{K}$, there exists a unique continuous map $f\colon Y\to\operatorname{Spc}^{cp}(\mathscr{K})$ such that $\sigma(x)=f^{-1}(\operatorname{supp}^{cp}(x))$. Explicitly, the map is defined by
@@
On the other hand, if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$, then $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp})$ is the only multiplicative support datum on $\mathscr{K}$, therefore it is trivially the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$.
@@
If every prime ideal of $\mathscr{K}$ is completely prime, then $\operatorname{Spc}^{cp}(\mathscr{K})\hookrightarrow\operatorname{Spc}(\mathscr{K})$ is the identity, hence $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ is multiplicative. Conversely, if $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ has the tensor product property, then the unique map map $f:\operatorname{Spc}(\mathscr{K})\to\operatorname{Spc}^{cp}(\mathscr{K})$ defined by $f(\mathscr{P})=\{a\in\mathscr{K}\mid\mathscr{P}\not\in\operatorname{supp}(a)\}$ is clearly the identity. Thus $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})=(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$, so $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ has the tensor product property. ∎
@@
Finally, we show $\operatorname{Spc}^{cp}$ is fully functorial, as promised.
@@
First we show well-definedness. It is easy to check that $F^{-1}(\mathscr{P})$ is a thick $\otimes$-ideal. For primality, suppose $a\otimes b\in F^{-1}(\mathscr{P})$, then $F(a)\otimes F(b)\in\mathscr{P}$. By primality, either $F(a)$ or $F(b)$ belongs to $\mathscr{P}$, hence $a$ or $b$ belongs to $F^{-1}(\mathscr{P})$ by definition. Note that this implies that if $\operatorname{Spc}^{cp}(\mathscr{L})$ is nonempty, then $\operatorname{Spc}^{cp}(\mathscr{K})$ necessarily is as well. On the other hand, if $\operatorname{Spc}^{cp}(\mathscr{L})$ is empty, this poses no problems since the empty set is the initial object in the category of topological spaces.
@@
$\operatorname{supp}^{cp}_{\mathscr{L}}(F(x))=\{\mathscr{Q}\in\operatorname{Spc}^{cp}(\mathscr{L})\mid F(x)\neq\mathscr{Q}\}.$
@@
$(\operatorname{Spc}F)^{-1}\big{(}\operatorname{supp}^{cp}_{\mathscr{K}}(x)\big{)}=\{\mathscr{Q}\in\operatorname{Spc}^{cp}(\mathscr{L})\mid x\not\in(\operatorname{Spc}(F))(\mathscr{Q})\}.$
@@
Since the compliments of supports form a basis of the topology of the complete prime spectrum, it follows that $\operatorname{Spc}(F)$ is continuous. Finally, $\operatorname{Spc}^{cp}(G\circ F)=\operatorname{Spc}^{cp}(F)\circ\operatorname{Spc}^{cp}(G)$ is straightforward. ∎
We make an easy observation: functoriality provides an easy check for when $\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$, the existence of a “fiber functor” to a more well-understood category with nonempty complete prime spectrum.
@@ 6. Radical thick tensor-ideals
Remark 6.3.
A pertinent question is whether every radical thick $\otimes$-ideal of $\mathscr{K}$ is an intersection of completely prime thick $\otimes$-ideals. At present, we are not aware of a counterexample.
On the other hand, an intersection of only prime thick $\otimes$-ideals is not necessarily radical, since not all prime thick $\otimes$-ideals are radical. In fact, for prime thick $\otimes$-ideals, being radical is equivalent to being completely prime.
Proposition 6.4.
Let $\mathscr{P}$ be a thick $\otimes$-ideal of $\mathscr{K}$. Then $\mathscr{P}$ is completely prime if and only if $\mathscr{P}$ is prime and radical.
Proof.
If $\mathscr{P}$ is completely prime, then $\mathscr{P}$ is prime and radical. Conversely, suppose for contradiction that $\mathscr{P}$ is prime and radical, but not completely prime. Then there exist $x,y\in\mathscr{K}$ for which $y\otimes x\in\mathscr{P}$ but neither $x\in\mathscr{P}$ nor $y\in\mathscr{P}$. Primality of $\mathscr{P}$ implies that there exists $z\in\mathscr{K}$ for which $x\otimes z\otimes y\not\in\mathscr{P}$. Then $\mathscr{P}$ radical implies that $(x\otimes z\otimes y)\otimes(x\otimes z\otimes y)\not\in\mathscr{P}$, but this is a contradiction since $y\otimes x\in\mathscr{P}$ and $\mathscr{P}$ is a $\otimes$-ideal. ∎